Suppose HTC sells 8,000 smartphones in Taiwan per week at a price of $ 250 . Due to competition

HTC reduces the price to $ 225 and subsequently sells 8640 units in the following week

b. Explain whether it was a good decision for HTC to reduce its price.


1
Expert's answer
2021-07-23T05:22:44-0400

(a) Let's first calculate the price elasticity of demand for smartphones:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""E_d=\dfrac{\dfrac{8640-8000}{0.5\cdot(8640+8000)}}{\dfrac{\$225-\$250}{0.5\cdot(\$225+\$250)}}=-0.73."

As we can see from calculations, "E_d<1". Therefore, the demand is inelastic.

(b) Since the demand is inelastic, it wasn't a good decision for HTC to reduce its price because the firm will lose its total revenue.